### v1

@@ 1 Introduction
Large Language Models (LLMs) such as GPT Achiam et al. (2023), LLaMA Touvron et al. (2023), OPT Zhang et al. (2022) and Falcon Institute (2023) have demonstrated strong performance across a wide range of advanced applications. However, serving LLMs is cost-demanding, requiring a large amount of hardware accelerators like GPUs to satisfy efficiency requirements such as latency and throughput.
Mainstream LLM serving systems primarily focus on high-performance GPUs like NVIDIA A100 and H100 in homogeneous GPU clusters. However, it is difficult for many LLM service providers to get access to sufficient high-performance GPUs, either due to the well known GPU shortage problem Yang et al. (2023), Strati et al. (2024) or the substantial fees. Meanwhile, with more and more advanced GPU architectures announced in the past few years, there are many less-performant GPUs in former generations remaining under-utilized. Thus, real-world cloud environments usually consist of heterogeneous GPUs and diverse prices. As shown in Table 1, cloud environments offer a wide range of hardware specifications and rental prices, providing users with diverse options to reduce the costs associated with LLM deployment and serving. Recent efforts Jiang et al. (2024), Mei et al. (2024), Griggs et al. (2024), Miao et al. (2023) have demonstrated that serving LLM with heterogeneous GPUs presents opportunities in reducing the serving cost. However, we find that these heterogeneous serving systems mainly address the heterogeneity in hardwares but fail to take account of the heterogeneity in the computation and memory-access workloads of different inference phases, which hinders the utilization of GPU resources. Such heterogeneity mainly comes from the distinct characteristics of LLM inference, and has raised a surge of research interests.
@@
3090Ti | 1008 GB/s | 40 TFLOPS | 24 GB | $0.307/hr |
Recent works Patel et al. (2023), Hu et al. (2024), Qin et al. (2024), Jin et al. (2024) have designed phase splitting approaches to utilize different amount of computational resources for prefill and decoding phase in LLM inference, which involves partitioning the two phases onto separate devices and transmitting the intermediate results (primarily KV caches) between them. Many empirical evidences have shown that such phase splitting approaches increase overall hardware utilization and system efficiency compared with the phase co-locating counterparts.
@@
Challenge 1: heterogeneous and limited resource pool. The available GPUs in cloud environments are usually in heterogeneous types, each with distinct specification (e.g., peak FLOPS, device memory bandwidth, and device memory limit), and the amount of each type is also restricted Yang et al. (2023), Strati et al. (2024). As a result, to deploy multiple copies (a.k.a. model replicas) of the same LLM, we must consider how to organize the available GPUs from a global view — given the available resources of diverse types, we need to jointly consider which GPUs should be grouped together to serve one model replica, and whether this replica should serve as the prefill or decoding phase. To our knowledge, this is an unexplored problem so far.
Challenge 2: heterogeneous and low network bandwidth. The second essential characteristic of cloud environments is that GPUs are usually connected through low network bandwidth, typically, PCIe for intra-node and ethernet for inter-node communication. And the network bandwidth also exhibits a high level of heterogeneity across different pairs of GPUs due to the discrepancy in connectivity (different PCIe versions, node locality, etc.). Such a network condition raises a hurdle for efficient LLM serving. On the one hand, transmitting KV caches from prefill to decode replicas inevitably incurs significant communication volume. While prior works Patel et al. (2023), Zhong et al. (2024) simply assume high-speed network connections (e.g., NVLink and Infini-Band) are available and overlook the communication overhead of KV caches (detailed in §2), which is impractical for clouds. On the other hand, due to the astonishing size of LLMs, model parallelism has been a cornerstone for LLM deployment. Thus, there expresses a need for designing heterogeneity-aware parallelization to facilitate phase splitting LLM serving on clouds.
Challenge 3: workload variability. Compared to in-house clusters, resources on clouds are more unstable Miao et al. (2023), Duan et al. (2024), Yousif (2018), Erben et al. (2024), and the distribution of requests (e.g., average arrival rate, input and output length) may change over time in online services in practice Wang et al. (2024a). These factors exacerbate the variability of serving workloads in cloud environments. In order to adapt to such workload variations, prior works Zhong et al. (2024) necessitate two steps: re-generating the deployment plan from scratch and re-loading the LLM parameters to adjust the model deployment. However, both steps are costly. Re-generating the deployment plan could take minutes to complete due to the complex hardware environments on cloud, and re-loading the LLM with a huge amount of parameters could be time-consuming. For instance, loading a 175B model with a disk bandwidth of 1.2 GBps takes over five minutes. Such expensive steps would lead to severe interruption to the online services.
@@ 2 Background and Related Works
Parallelism strategies. To parallelize the model over multiple GPUs, there are two prevalent forms of model parallelism, which are tensor model parallelism and pipeline model parallelism. Tensor model parallelism (TP) Shoeybi et al. (2019), Nagrecha (2021) divides model weights and computationally intensive operations such as matrix multiplication across various GPUs, thereby splitting data scanning and computation to minimize LLM inference latency, particularly the TTFT in the prefill phase. Pipeline model parallelism (PP) Huang et al. (2019), Narayanan et al. (2019) divides the layers of a model into multiple stages. These stages are assigned to distinct GPUs for execution and they establish a pipeline. Only inter-layer activations are needed to be communicated between stages.
Phase splitting deployment. As the prefill and decode phases differ in workload characteristics (i.e., compute/memory-bound) significantly, recent efforts propose to utilize different hardware resources for the two phases in order to avoid performance interference Patel et al. (2023), Zhong et al. (2024), Hu et al. (2024), Jin et al. (2024), Qin et al. (2024). In other words, there are two kinds of model replicas, one for prefill and the other for decode, respectively, and it is necessary to transmit the KV cache from the prefill replicas to the decode ones. Due to the substantial size of KV cache, existing efforts essentially require high communication bandwidth for the KV cache transfer — Zhong et al. (2024) proposed to colocate prefill and decode replicas on GPUs within the same node, facilitating fast KV cache transfer with NVLink, while Patel et al. (2023) and Hu et al. (2024) utilized high-speed Infini-Band connections for inter-node communication. However, in cloud environments, GPUs are connected via limited bandwidth (typically, PCIe for intra-node connections and Ethernet for inter-node connections) rather than high-speed connections (typically, NVLink and Infini-Band). Therefore, KV cache transfer would lead to a huge cost and it necessitates enhancement.
Heterogeneous GPU Computing. Recent research has investigated diverse approaches to deploy large models on heterogeneous GPU clusters. HexGen Jiang et al. (2024) proposes asymmetric partitioning and advanced scheduling to deploy generative inference in a decentralized and heterogeneous setting. Helix Mei et al. (2024) formulates the heterogeneous GPUs and network connections as a maxflow problem and adopts a mixed integer linear programming algorithm to discover highly optimized strategies to serve LLMs. Our work has a similar objective, but is the first effort that integrates phase splitting with the heterogeneous GPUs to provide high-performance cloud serving for LLMs.
@@ 3.2 Solving the Upper-level Problem
In cloud environments, there are various types of GPUs with heterogeneous capabilities, making the group construction a non-trivial problem. In addition, the phase-splitting design requires meticulous phase designation to the model serving groups for better performance, making the upper-level problem even more complex. Formally, in Appendix A, we show that the upper-level problem is essentially a job shop scheduling problem (JSSP), which is a notoriously difficult NP-hard problem in combinatorial optimization Sotskov & Shakhlevich (1995), Omar et al. (2006).
A well-known approach to solve JSSP is tabu search Glover (1990), Glover & Laguna (1998), Gendreau & Potvin (2005) and there have been many efforts applying tabu search to solve JSSP under various situations Hurink et al. (1994), Dauzère-Pérès & Paulli (1997), Zhang et al. (2007). Motivated by this, we adapt tabu search to the upper-level problem and design a brand new algorithm to identify the optimal deployment plan, which is demonstrated in Algorithm 1. In essence, it starts from an initial solution, and leverages an iterative neighborhood search process to improve the solution. Below we focus on how to determine the initial solution and how to construct neighbors given the current solution in our scenario.
Initialization. It is essential to have a good initial solution in tabu search in order to speedup the search process and escape from local optima. Thus, we utilize the Hierarchical Clustering method Shetty & Singh (2021) to cluster the GPUs according to their inter-connection bandwidth matrix, and subsequently treat each generated cluster as one model serving group at initialization. Intuitively, this makes the initial assignment of model serving groups strategically avoid connections with ultra-low communication bandwidth in the cloud environment. In addition, the phase designation of each group is randomly initialized.
@@ 3.3 Solving the Lower-level Problem
Numerous studies Jiang et al. (2024), Zheng et al. (2022), Li et al. (2023), Miao et al. (2022; 2023), Wang et al. (2024b) have investigated how to deduce the optimal parallel configuration by meticulously enumerating a vast number of possible configurations. In this work, we further take the key characteristics of cloud environments into account and design heuristics to accelerate the enumeration process. We introduce the heuristics below and leave the details of the deduction routine in Appendix B.
@@
We adopt the inference task simulator from DistServe Zhong et al. (2024), which estimates the SLO attainment according to workload information (e.g., input length, output length, etc.) and single request processing time. It is noteworthy that we integrate the KV cache communication cost into the simulator, since it is non-negligible in cloud environments. To be specific, suppose there are prefill replicas and decode replicas, our simulator enumerates every pair of them and estimates the SLO attainment by integrating the KV cache communication cost, which is analyzed via the alpha-beta model Hockney (1994):
@@
Based on this, we estimate the SLO attainment of every pair of prefill and decode replicas. Formally, denote as the SLO attainment matrix, where represents the estimated SLO attainment when requests are processed by the -th prefill replica and the -th decode replica. Then, we turn the optimization problem of overall system SLO attainment into a simple two-stage transportation problem (TSTP) Santoso & Heryanto (2022) as follows:
@@ 3.4 Lightweight Rescheduling in Real Time
Numerous factors in cloud services affect the optimal deployment plan, with two primary factors being LLM inference workloads and GPU availability. On the one hand, LLM services usually exhibit significant variation in workload characteristics across different downstream tasks. For instance, coding workloads typically generate shorter responses than conversational workloads but usually have longer prompts Patel et al. (2023). On the other hand, compared to in-house clusters, cloud resources are inherently more dynamic and unstable, necessitating a good support for cluster size adjustments in real-time. Consequently, rescheduling is essential for ThunderServe to adapt the deployment plan to varying online workloads and cluster size changes on cloud. However, altering the deployment plan is far from trivial in cloud environments. If we re-run the scheduling algorithm from scratch and reload the model parameters according to the updated deployment plan (take minutes to complete), it would lead to severe interruption to the online services. Therefore, we propose a lightweight rescheduling process that only adjusts the phase designation and orchestration in the deployment plan to accommodate varying workloads and cluster sizes.
@@ 4 Implementation
ThunderServe is a distributed LLM serving system designed to optimize online services in cloud environments, which develops a novel scheduling algorithm to partition the given cloud GPU resources into model serving groups, designate which phase each group should serve as, deduce the optimal parallel configuration for each group, and determine how the requests should be routed among groups. It is implemented using 20K lines of Python and C++/CUDA code. Besides, ThunderServe incorporates FlashAttention Dao et al. (2022) and PagedAttention Kwon et al. (2023) to accelerate LLM inference, and leverages the batching strategy proposed by Zhong et al. (2024) for LLM serving.
@@
KV cache compression technique. As discussed in §2, prior works rely on high-bandwidth connections (i.e., NVLINK or InfiniBand) for transferring KV cache in phase splitting deployment, which is impractical in cloud service scenarios characterized by heterogeneous network conditions among GPUs. To reduce the KV cache communication cost, we borrow the idea of low-precision quantization from KIVI Liu et al. (2024) to quantize KV cache to fewer bits, so that the size of each element (i.e. in Equation 1) is shrinked. However, unlike existing works in the field of KV cache quantization Liu et al. (2024), Kang et al. (2024), our system does not retain low bitwidths when using the KV cache values for computation. Specifically, the KV cache values in the prefill replica are quantized and packed for communication, and then immediately unpacked and dequantized after they are received by the decode replica. Thus, both the prefill and decode phases are conducted using the 16-bit KV cache values rather than the quantized ones. By this means, we can significantly reduce the KV cache communication volume, without harming the model quality.
@@ 5.1 Experimental Setup
Model and workloads. We deploy the popular open-source LLaMA-30B model across two real-world workloads, coding and conversation, from the Azure Conversation dataset Patel et al. (2023). And we follow prior works Li et al. (2023), Jiang et al. (2024) to generate the inference workload using a Poisson process determined by the request rate, with consecutive requests (inter-arrival times) following an exponential distribution.
Evaluation metrics. Following prior works Patel et al. (2023), Zhong et al. (2024), we focus on overall system SLO attainment and throughput when evaluating the performance. System SLO attainment indicates the percentage of requests completed within a predefined latency deadline. There are three types of SLO: TTFT, TPOT, and E2E SLO. We specifically measure system SLO attainment by the percentage of requests that meet the time criteria established by each SLO type. We scale the SLO to various multiples of the execution latency of A100 GPUs (SLO Scale in Figure 8), which allows us to evaluate system performance under different levels of operational stringency. For a target SLO attainment goal (e.g., 90% and 99%), we focus on the minimum latency deadline required to achieve the desired attainment.
Baselines. We consider state-of-the-art systems under both the cloud and in-house settings, respectively. Under the cloud setting, we consider HexGen Jiang et al. (2024), an LLM serving system for clouds featuring advanced scheduling and asymmetric parallelism. For the in-house scenario, we consider vLLM Kwon et al. (2023), a prestigious LLM serving system, as well as DistServe Zhong et al. (2024), an LLM serving system featuring phase splitting.

### v2

@@ 1 Introduction
Large Language Models (LLMs) such as GPT achiam2023gpt, LLaMA touvron2023llama, OPT zhang2022opt and Falcon falcon180b have demonstrated strong performance across a wide range of advanced applications. However, serving LLMs is cost-demanding, requiring a large amount of hardware accelerators like GPUs to satisfy efficiency requirements such as latency and throughput.
Mainstream LLM serving systems primarily focus on high-performance GPUs like NVIDIA A100 and H100 in homogeneous GPU clusters. However, it is difficult for many LLM service providers to get access to sufficient high-performance GPUs, either due to the well known GPU shortage problem skypilot, euromlsysworkshop or the substantial fees. Meanwhile, with more and more advanced GPU architectures announced in the past few years, there are many less-performant GPUs in former generations remaining under-utilized. Thus, real-world cloud environments usually consist of heterogeneous GPUs and diverse prices. As shown in Table 1, cloud environments offer a wide range of hardware specifications and rental prices, providing users with diverse options to reduce the costs associated with LLM deployment and serving. Recent efforts jiang2024hexgen, mei2024helix, griggs2024melange, miao2023spotserve have demonstrated that serving LLM with heterogeneous GPUs presents opportunities in reducing the serving cost. However, we find that these heterogeneous serving systems mainly address the heterogeneity in hardwares but fail to take account of the heterogeneity in the computation and memory-access workloads of different inference phases, which hinders the utilization of GPU resources. Such heterogeneity mainly comes from the distinct characteristics of LLM inference, and has raised a surge of research interests.
@@
| 3090Ti | 1008 GB/s | 71 TFLOPS | 24 GB | $0.307/hr |
Recent works patel2023splitwise, hu2024inference, qin2024mooncake, jin2024p have designed phase splitting approaches to utilize different amount of computational resources for prefill and decoding phase in LLM inference, which involves partitioning the two phases onto separate devices and transmitting the intermediate results (primarily KV caches) between them. Many empirical evidences have shown that such phase splitting approaches increase overall hardware utilization and system efficiency compared with the phase co-locating counterparts.
@@
Challenge 1: heterogeneous and limited resource pool. The available GPUs in cloud environments are usually in heterogeneous types, each with distinct specification (e.g., peak FLOPS, device memory bandwidth, and device memory limit), and the amount of each type is also restricted skypilot, euromlsysworkshop. As a result, to deploy multiple copies (a.k.a. model replicas) of the same LLM, we must consider how to organize the available GPUs from a global view — given the available resources of diverse types, we need to jointly consider which GPUs should be grouped together to serve one model replica, and whether this replica should serve as the prefill or decoding phase. To our knowledge, this is an unexplored problem so far.
Challenge 2: heterogeneous and low network bandwidth. The second essential characteristic of cloud environments is that GPUs are usually connected through low network bandwidth, typically, PCIe for intra-node and ethernet for inter-node communication. And the network bandwidth also exhibits a high level of heterogeneity across different pairs of GPUs due to the discrepancy in connectivity (different PCIe versions, node locality, etc.). Such a network condition raises a hurdle for efficient LLM serving. On the one hand, transmitting KV caches from prefill to decode replicas inevitably incurs significant communication volume. While prior works patel2023splitwise, zhong2024distserve simply assume high-speed network connections (e.g., NVLink and Infini-Band) are available and overlook the communication overhead of KV caches (detailed in §2), which is impractical for clouds. On the other hand, due to the astonishing size of LLMs, model parallelism has been a cornerstone for LLM deployment. Thus, there expresses a need for designing heterogeneity-aware parallelization to facilitate phase splitting LLM serving on clouds.
Challenge 3: workload variability. Compared to in-house clusters, resources on clouds are more unstable miao2023spotserve, duan2024parcae, yousif2018cloud, erben2024can, and the distribution of requests (e.g., average arrival rate, input and output length) may change over time in online services in practice wang2024burstgpt. These factors exacerbate the variability of serving workloads in cloud environments. In order to adapt to such workload variations, prior works zhong2024distserve necessitate two steps: re-generating the deployment plan from scratch and re-loading the LLM parameters to adjust the model deployment. However, both steps are costly. Re-generating the deployment plan could take minutes to complete due to the complex hardware environments on cloud, and re-loading the LLM with a huge amount of parameters could be time-consuming. For instance, loading a 175B model with a disk bandwidth of 1.2 GBps takes over five minutes. Such expensive steps would lead to severe interruption to the online services.
@@ 2 Background and Related Works
Parallelism strategies. To parallelize the model over multiple GPUs, there are two prevalent forms of model parallelism, which are tensor model parallelism and pipeline model parallelism. Tensor model parallelism (TP) shoeybi2019megatron, nagrecha2021model divides model weights and computationally intensive operations such as matrix multiplication across various GPUs, thereby splitting data scanning and computation to minimize LLM inference latency, particularly the TTFT in the prefill phase. Pipeline model parallelism (PP) huang2019gpipe, narayanan2019pipedream divides the layers of a model into multiple stages. These stages are assigned to distinct GPUs for execution and they establish a pipeline. Only inter-layer activations are needed to be communicated between stages.
Phase splitting deployment. As the prefill and decode phases differ in workload characteristics (i.e., compute/memory-bound) significantly, recent efforts propose to utilize different hardware resources for the two phases in order to avoid performance interference patel2023splitwise, zhong2024distserve, hu2024inference, jin2024p, qin2024mooncake. In other words, there are two kinds of model replicas, one for prefill and the other for decode, respectively, and it is necessary to transmit the KV cache from the prefill replicas to the decode ones. Due to the substantial size of KV cache, existing efforts essentially require high communication bandwidth for the KV cache transfer — zhong2024distserve proposed to colocate prefill and decode replicas on GPUs within the same node, facilitating fast KV cache transfer with NVLink, while patel2023splitwise and hu2024inference utilized high-speed Infini-Band connections for inter-node communication. However, in cloud environments, GPUs are connected via limited bandwidth (typically, PCIe for intra-node connections and Ethernet for inter-node connections) rather than high-speed connections (typically, NVLink and Infini-Band). Therefore, KV cache transfer would lead to a huge cost and it necessitates enhancement.
Heterogeneous GPU Computing. Recent research has investigated diverse approaches to deploy large models on heterogeneous GPU clusters. HexGen jiang2024hexgen proposes asymmetric partitioning and advanced scheduling to deploy generative inference in a decentralized and heterogeneous setting. Helix mei2024helix formulates the heterogeneous GPUs and network connections as a maxflow problem and adopts a mixed integer linear programming algorithm to discover highly optimized strategies to serve LLMs. Our work has a similar objective, but is the first effort that integrates phase splitting with the heterogeneous GPUs to provide high-performance cloud serving for LLMs.
@@ 3.2 Solving the Upper-level Problem
In cloud environments, there are various types of GPUs with heterogeneous capabilities, making the group construction a non-trivial problem. In addition, the phase-splitting design requires meticulous phase designation to the model serving groups for better performance, making the upper-level problem even more complex. Formally, in Appendix A, we show that the upper-level problem is essentially a job shop scheduling problem (JSSP), which is a notoriously difficult NP-hard problem in combinatorial optimization sotskov1995np, omar2006job.
A well-known approach to solve JSSP is tabu search glover1990tabu, glover1998tabu, gendreau2005tabu and there have been many efforts applying tabu search to solve JSSP under various situations hurink1994tabu, tabu_jssp_1997, zhang2007tabu. Motivated by this, we adapt tabu search to the upper-level problem and design a brand new algorithm to identify the optimal deployment plan, which is demonstrated in Algorithm 1. In essence, it starts from an initial solution, and leverages an iterative neighborhood search process to improve the solution. Below we focus on how to determine the initial solution and how to construct neighbors given the current solution in our scenario.
Initialization. It is essential to have a good initial solution in tabu search in order to speedup the search process and escape from local optima. Thus, we utilize the Hierarchical Clustering method shetty2021hierarchical to cluster the GPUs according to their inter-connection bandwidth matrix, and subsequently treat each generated cluster as one model serving group at initialization. Intuitively, this makes the initial assignment of model serving groups strategically avoid connections with ultra-low communication bandwidth in the cloud environment. In addition, the phase designation of each group is randomly initialized.
@@ 3.3 Solving the Lower-level Problem
Numerous studies jiang2024hexgen, zheng2022alpa, li2023alpaserve, Miao_2022, miao2023spotserve, wang2024improving have investigated how to deduce the optimal parallel configuration by meticulously enumerating a vast number of possible configurations. In this work, we further take the key characteristics of cloud environments into account and design heuristics to accelerate the enumeration process. We introduce the heuristics below and leave the details of the deduction routine in Appendix B.
@@
We adopt the inference task simulator from DistServe zhong2024distserve, which estimates the SLO attainment according to workload information (e.g., input length, output length, etc.) and single request processing time. It is noteworthy that we integrate the KV cache communication cost into the simulator, since it is non-negligible in cloud environments. To be specific, suppose there are prefill replicas and decode replicas, our simulator enumerates every pair of them and estimates the SLO attainment by integrating the KV cache communication cost, which is analyzed via the alpha-beta model hockney1994communication:
@@
Based on this, we estimate the SLO attainment of every pair of prefill and decode replicas. Formally, denote as the SLO attainment matrix, where represents the estimated SLO attainment when requests are processed by the -th prefill replica and the -th decode replica. Then, we turn the optimization problem of overall system SLO attainment into a simple two-stage transportation problem (TSTP) santoso2022development as follows:
@@ 3.4 Lightweight Rescheduling in Real Time
Numerous factors in cloud services affect the optimal deployment plan, with two primary factors being LLM inference workloads and GPU availability. On the one hand, LLM services usually exhibit significant variation in workload characteristics across different downstream tasks. For instance, coding workloads typically generate shorter responses than conversational workloads but usually have longer prompts patel2023splitwise. On the other hand, compared to in-house clusters, cloud resources are inherently more dynamic and unstable, necessitating a good support for cluster size adjustments in real-time. Consequently, rescheduling is essential for ThunderServe to adapt the deployment plan to varying online workloads and cluster size changes on cloud. However, altering the deployment plan is far from trivial in cloud environments. If we re-run the scheduling algorithm from scratch and reload the model parameters according to the updated deployment plan (take minutes to complete), it would lead to severe interruption to the online services. Therefore, we propose a lightweight rescheduling process that only adjusts the phase designation and orchestration in the deployment plan to accommodate varying workloads and cluster sizes.
@@ 4 Implementation
ThunderServe is a distributed LLM serving system designed to optimize online services in cloud environments, which develops a novel scheduling algorithm to partition the given cloud GPU resources into model serving groups, designate which phase each group should serve as, deduce the optimal parallel configuration for each group, and determine how the requests should be routed among groups. It is implemented using 20K lines of Python and C++/CUDA code. Besides, ThunderServe incorporates FlashAttention dao2022flashattention and PagedAttention kwon2023efficient to accelerate LLM inference, and leverages the batching strategy proposed by zhong2024distserve for LLM serving.
@@
KV cache compression technique. As discussed in §2, prior works rely on high-bandwidth connections (i.e., NVLINK or InfiniBand) for transferring KV cache in phase splitting deployment, which is impractical in cloud service scenarios characterized by heterogeneous network conditions among GPUs. To reduce the KV cache communication cost, we borrow the idea of low-precision quantization from KIVI kivi2024 to quantize KV cache to fewer bits, so that the size of each element (i.e. in Equation 1) is shrinked. However, unlike existing works in the field of KV cache quantization kivi2024, kang2024gear, our system does not retain low bitwidths when using the KV cache values for computation. Specifically, the KV cache values in the prefill replica are quantized and packed for communication, and then immediately unpacked and dequantized after they are received by the decode replica. Thus, both the prefill and decode phases are conducted using the 16-bit KV cache values rather than the quantized ones. By this means, we can significantly reduce the KV cache communication volume, without harming the model quality.
@@ 5.1 Experimental Setup
Model and workloads. We deploy the popular open-source LLaMA-30B model across two real-world workloads, coding and conversation, from the Azure Conversation dataset patel2023splitwise. And we follow prior works li2023alpaserve, jiang2024hexgen to generate the inference workload using a Poisson process determined by the request rate, with consecutive requests (inter-arrival times) following an exponential distribution.
Evaluation metrics. Following prior works patel2023splitwise, zhong2024distserve, we focus on overall system SLO attainment and throughput when evaluating the performance. System SLO attainment indicates the percentage of requests completed within a predefined latency deadline. There are three types of SLO: TTFT, TPOT, and E2E SLO. We specifically measure system SLO attainment by the percentage of requests that meet the time criteria established by each SLO type. We scale the SLO to various multiples of the execution latency of A100 GPUs (SLO Scale in Figure 8), which allows us to evaluate system performance under different levels of operational stringency. For a target SLO attainment goal (e.g., 90% and 99%), we focus on the minimum latency deadline required to achieve the desired attainment.
Baselines. We consider state-of-the-art systems under both the cloud and in-house settings, respectively. Under the cloud setting, we consider HexGen jiang2024hexgen, an LLM serving system for clouds featuring advanced scheduling and asymmetric parallelism. For the in-house scenario, we consider vLLM kwon2023efficient, a prestigious LLM serving system, as well as DistServe zhong2024distserve, an LLM serving system featuring phase splitting.